OpenStudy (anonymous):

Last year, there were 120 seniors in Limon High School. This year there are 145. What is the percent of increase? Round to the nearest percent. A. 83% B. 17% C. 21% D. 79%

OpenStudy (campbell_st):

well is simply the increase in the number of students/previous number x 100 or \[\frac{145 -120}{120} \times 100 =\]

OpenStudy (anonymous):

21%?

OpenStudy (campbell_st):

thats it